OpenStudy (anonymous):

A direct variation function contains the points (2, 14) and (4, 28). Which equation represents the function? y = x/14 y = x/7 y = 7x y = 14x

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

2 times what is 14 ?

OpenStudy (danjs):

The general way the books tell you to solve these, is to use the direct variation thing y = k*x and use your points (x,y) in there to figure out what k is.

OpenStudy (anonymous):

@dumbcow 7

OpenStudy (danjs):

(x,y) = (2,14), x=2 and y=14 Put that into y = k*x

OpenStudy (danjs):

14 = k*2 k = 7 y = 7x

OpenStudy (dumbcow):

there you go ..... its multiplying by 7

OpenStudy (danjs):

That is the basic way to solve the direct variation problems, use y = k*x put a given point in, and find out what 'k' is